Lucky Buy Company’s stock has suffered due to several warranty-related lawsuits filed against the...

  1. Lucky Buy Company’s stock has suffered due to several warranty-related lawsuits filed against the company. The company revamped their warranty program and introduced additional customer services to improve the customer experience. Yet, its stock price is only $13 per share. Management is planning a two-for-seven reverse stock split to increase the stock price and bring it closer to the average stock price in the industry. Assume that John Thornton, the last remaining co-founder of the firm, owns 420,000 shares. (7 points)
  1.    How many shares will he own after the reverse stock split?
  2. What is the anticipated price of the stock after the reverse stock split?
  3.    Because investors often have a negative reaction to a reverse stock split, assume the stock only goes up to 85 percent of the value computed in part b. What will the stock’s price be?
  4. How will the total value of John Thornton’s holdings change from before the reverse stock split to after the reverse stock split (based on the stock value computed in part c)?
Answer #1

A reverse split is one in which the company reduces its number of shares outstanding to boost its stock price .
a) For a reverse stock split of two for seven, the company will take away seven stocks and give back two stocks.
Therefore, if John has 420,000 shares then after stock split he will own
= 420000 / (7/2)
= 120,000

He will own 120,000 shares after stock split.

b) Market cap of the stock would remain the same even after reverse stock split.
Before reverse stock split ,
Market cap = 420000 * 13

After reverse stock split
Market cap = 120000 * X
where X is the new anticipated price of the stock

420000 * 13 = 120000 * X
X = 13 * 420000 / 120000
   = $45.5

Anticipated price would be $45.5


c) 85% *45.5
= $38.675

d) Before reverse stock split the total value was = 420000 * 13 = $5460000

after reverse stock split the total value = 120000 * 38.675 = $4641000

Therefore, the total value has declined by ( $5460000 - $4641000 )= $819000

% decline = 819000 / 5460000
                 = 15%